OpenStudy (anonymous):

AHRIGHT NOW KIDDOS HOW DO U GIT THE SQUARE ROOT OF 18 HELP A BRUDDA OUT

OpenStudy (btaylor):

If you want a decimal value, use a calculator. Otherwise, factor the number, and pull out any pairs. \[\sqrt{18} = \sqrt{2 \times \boxed{3 \times 3}} \rightarrow 3 \sqrt{2}\]

OpenStudy (anonymous):

does it help to know that \(18=9\times 2\)?

OpenStudy (solomonzelman):

what in the world is your question?

OpenStudy (solomonzelman):

Ask your teacher in class like that I dare.

OpenStudy (anonymous):

Take your calculator and do \( \sqrt{18} \) or you could see if there is a perfect square and a non perfect square and there is. \( \sqrt{9} \times \sqrt{2} = 3\sqrt{2}\)

OpenStudy (anonymous):

Yeah I was a little slow to the party lol

OpenStudy (anonymous):

shut up soloman zelman aint nobody axt u 4 ur opinionz

pooja195 (pooja195):

Yo learn to be nice you are on a study site if you are simply trying to waste peoples time close this.

OpenStudy (anonymous):

When working with squares you can bust up numbers into perfect and non perfect squares. This is not always the case but sometimes you can and in your situations you could. A perfect square is a number that can be squared evenly . For instance, 4 is a perfect square. \( \sqrt{4} = 2 \) a non perfect square is like the following \( \sqrt{2} \)

OpenStudy (anonymous):

umm excuse me pooja who in the hell told you to drag yaself into this

OpenStudy (anonymous):

Leah that was totally uncalled for

OpenStudy (anonymous):

watch ur mouth, @pooja195 is an ambassador and lol yah I know

OpenStudy (anonymous):

its okay leahtoham i heart u

pooja195 (pooja195):

\(\huge\color{orange}{Welcome~To~OpenStudy}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

pooja195 (pooja195):

Please do look at the CoC again.

OpenStudy (anonymous):

You tell em!

pooja195 (pooja195):

http://openstudy.com/code-of-conduct